OpenStudy (anonymous):

4/9 (x+3)=g Solve for X. State any restrictions on the variables

OpenStudy (anonymous):

is this a multiple Q:

OpenStudy (anonymous):

Yes 4/9 is multiplied to (x+3)

OpenStudy (anonymous):

4/9(x+3)=g Take off the 4/9 by multiplying by the inverse on both sides of the equation: (9/4)*(4/9)(x+3)=g*(9/4) That'll cancel out leaving you with this: x+3=(9/4)g Subtract the 3 from both sides of the equation: X+3-3=(9/4)g-3 That leaves you with X: X=((9/4)g)-3